OpenStudy (anonymous):

Hello! Not quite sure about how to find derivative of: y = 3x - 5cos('pi'x)^2

OpenStudy (anonymous):

3x prime=3, -5cos(u)^2= 10sin(u)*u prime= 3+10sin(pi*x)*pi

OpenStudy (anonymous):

y'=3-5*2*cos(pi*x)*pi*(-sin(pi*x))=3+5pi*sin(2pix)

OpenStudy (anonymous):

dellcom, you forgot that (cos^2)'=2cos*(-sin).... :)

OpenStudy (anonymous):

OK, so I use the chain use. But I'm still missing something. The answer I get is: \[y = 3 + 10\cos \pi x \sin \pi x\] Its seems like you combined something that I am not seeing..